# Measuring deviations from spherical symmetry

### v1

@@ Abstract
Most of the work on checking spherical symmetry assumptions on the distribution of $p$-dimensional random vector $Y$ has its focus on statistical tests for the null hypothesis of exact spherical symmetry. In this paper we take a different point of view and propose a measure for the deviation from spherical symmetry, which is based on the minimum distance between the distribution of the vector $\big(\|Y\|,Y/\|Y\|)^{\top}$ and its best approximation by a distribution of a vector $\big(\|Y_{s}\|,Y_{s}/\|Y_{s}\|)^{\top}$ corresponding to a random vector $Y_{s}$ with a spherical distribution. We develop estimators for the minimum distance with corresponding statistical guarantees (provided by asymptotic theory) and demonstrate the applicability of our approach by means of a simulation study and a real data example.
@@ 1 Introduction
Spherical contoured models include multivariate Gaussian distribution, multivariate $t$- distribution and multivariate stable distribution find their applications in earth science, vector cardiology (Mardia and Jupp,, 1999; Di Marzio et al.,, 2019), portfolio theory (Gupta et al.,, 2013), Mahalanobis distances (Luis Angel García-Escudero,, 2005). There has been a wide interest in testing for the spherical symmetry, or a wider class of elliptical symmetry, see for example Smith, (1977), Baringhaus, (1991), Fang et al., (1993), Koltchinskii and Li, (1998), Schott, (2002), Huffer and Park, (2007), Liang et al., (2008), Cassart et al., (2008), Henze et al., (2014), Albisetti et al., (2020), Babić et al., (2021), Banerjee and Ghosh, (2024) among others. A common feature of most of the cited references is that the authors use different approaches to construct tests for the null hypothesis of exact spherical symmetry. However, this assumption is rarely met because one believes that it is exactly (with mathematical equality) satisfied, but with the hope that the deviation from the null hypothesis is small such that (optimal) inference under the assumption of spherical symmetry is still reliable. Thus, strictly speaking, there are many situations where tests for exact sphericity are performed, although it is clear that this hypothesis is at most “approximately” satisfied.
@@
In this paper, we take a different point of view on the problem of validating the assumption of spherical symmetry on the distribution of a random vector $Y$. Instead of testing for spherical symmetry, we propose a measure of deviation from spherical symmetry. This measure is based on the minimum distance between the distribution of the vector $\big(\|Y\|,Y/\|Y\|)^{\top}$ and its best approximation by a distribution of a vector $\big(\|Y_{s}\|,Y_{s}/\|Y_{s}\|)^{\top}$ corresponding to a vector $Y_{s}$ with a spherical distribution, where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{p}$ (note that a random variable $Y_{s}$ has a spherical distribution if and only if $\|Y_{s}\|$ and $Y_{s}/\|Y_{s}\|$ are independent and $Y_{s}/\|Y_{s}\|$ is uniformly distributed on the $p$-dimensional sphere $\mathbb{S}_{p-1}$). In Section 2 we derive an explicit expression for this minimum distance, which is essentially the difference between the (squared) $L^{2}$-norm of the density of the vector $\big(\|Y_{s}\|,Y_{s}/\|Y_{s}\|)^{\top}$ and the (squared) $L^{2}$-norm of the density of the random variable $\|Y_{s}\|$ multiplied with the surface area of $\mathbb{S}^{p-1}$. Consequently, the estimation problem of the minimum distance boils down to estimating the squared $L^{2}$-norm of the density of a distribution on the sphere and a distribution on $\mathbb{R}^{+}$. While the problem of estimating the integrated squared density on $\mathbb{R}^{p}$ has been extensively studied from various perspectives including optimality and adaptivity (see Hall and Marron,, 1987; Bickel and Ritov,, 1988; Laurent and Massart,, 2000; Giné and Nickl,, 2008, among many others), the corresponding problem on the sphere have not found much attention so far. We construct an estimator for the minimum distance from a sequence of independent identically distributed observations combining an estimator for the integrated squared density of the joint distribution of $Y/\|Y\|$ and $\|Y\|$ with an estimator for the integrated squared density of $\|Y\|$ on the non-negative line. We prove asymptotic normality of an appropriately centered and normalized version. Interestingly, from the theoretical perspective, there appear to be different scalings, depending on whether the distribution is spherical symmetric or not. In the latter case, the difference of the estimators converges at a rate $\sqrt{n}$ as in the case estimating the integral of the squared density (see, for example, Bickel and Ritov,, 1988). On the other hand, if the distribution of $Y$ is in fact spherical symmetric, the estimator converges at a rate $\sqrt{n(n-1)h/{\kappa^{(p-1)/2}}}$, where $h\to 0$ and $\kappa\to\infty$ denote smoothing parameters required for the estimation of the integrated squared densities of $\|Y\|$ and $Y/\|Y\|$, respectively.
@@
Based on these results, we develop several statistical applications. In Section 3.2 we derive an asymptotic confidence interval for the deviation of the distribution from sphericity. Moreover, we also construct tests for the hypothesis that the deviation from sphericity exceeds a given threshold. For example, if ${\cal M}^{2}$ denotes the measure of deviation we derive a consistent and asymptotic level $\alpha$-test for the hypotheses $H_{0}^{\rm eq}:{\cal M}^{2}>\Delta$ versus $H_{1}^{\rm eq}:{\cal M}^{2}\leq\Delta$, which allows to decide for approximate sphericity at a controlled type $I$ error. These procedures require the estimation of a complicated asymptotic variance, which we address by a jackknife procedure. Moreover, as an alternative, we also develop in Section 4 pivotal inference for the measure ${\cal M}^{2}$, which does not require any estimation of asymptotic variances. Our approach is based on the self-normalization principle (see Shao,, 2015), which cannot be used directly in the present context, but has to be adapted to the specific statistical problems under consideration. In particular, we have to prove the weak convergence of a sequential estimator of ${\cal M}^{2}$ to obtain a pivotal statistic which can be used for the construction of confidence intervals and for testing the hypotheses of a relevant deviation from sphericity. Finally, we demonstrate the applicability of our approach through a small simulation study and a real data example.
@@ 2 A measure of deviation and its estimation
for some function $g$. Consequently, observing (2.1), it follows that $Y$ has a spherical symmetric distribution if and only $U=\|Y\|$ and $V=Y/\|Y\|$ are independent and $V=Y/\|Y\|$ is uniformly distributed on the sphere with (constant) density
@@
Let $Y_{1},\ldots,Y_{n}$ denote independent identically distributed $p$-dimensional random variables and define $U_{i}=\|Y_{i}\|$; $V_{i}=Y_{i}/\|Y_{i}\|$ ($i=1,\ldots,n)$. By Proposition 2.1 the problem of estimating the minimum distance ${\cal M}^{2}$ boils down to the estimation of the integrated squared densities of the random variables $(U_{i},V_{i})$ and $U_{i}$. The latter can be estimated by standard methods (see, for example Hall and Marron,, 1987; Bickel and Ritov,, 1988, among many others) and we use
@@
as estimate for the integrated squared density of $U$, where $h$ denotes a bandwidth. The estimation of the first integral in (2.5) is more intricate as it corresponds to the (squared) $L^{2}$-norm of a density on $\mathbb{R}^{+}\times\mathbb{S}^{p-1}$. For this purpose, we use classical results from density estimation on the sphere, such as Hall et al., (1987), and consider a rapidly varying function $L:\mathbb{R}^{p}\to\mathbb{R}$, which satisfies (i) $L$ is nonnegative and nondecreasing, (ii) for each $0<r<1$,
@@
$\hat{\cal M}_{n}^{(2)}:=\frac{1}{n(n-1)c_{1}(\kappa)h}\sum_{i\neq j}K\left(\frac{U_{i}-U_{j}}{h}\right)L(\kappa V_{i}^{\top}V_{j})$ (2.8)
where $K$ is a kernel function and $h$ is the same bandwidth as used in the estimator $\hat{\cal M}_{n}^{(1)}$, $\kappa\to\infty$ is another bandwidth corresponding to the spherical part of the vector $(U,V)$ and the constant $c_{1}(\kappa)$ is given by
@@
for the kernel $L$, where $\mathcal{I}_{\nu}$ is the modified Bessel function of the first kind (see also 9.6.18 on Page 376 of Abramowitz and Stegun,, 1968) of order $\nu$, i.e.,
@@ 3.1 Asymptotic theory
For the investigation of the bias of $\hat{\mathcal{M}}_{n}^{2}$, we recall a Taylor expansion for a twice differentiable function $f:\mathbb{R}^{+}\times\mathbb{S}^{p-1}\to\mathbb{R}$ with respect to the argument $v$ on the sphere $\mathbb{S}^{p-1}$. According to equation (3) in Marco Di Marzio and Taylor, (2014), we have
$f(u,y)=f(u,x)+\theta\xi^{\top}\mathcal{D}_{f}(u,x)+\frac{\theta^{2}}{2}\xi^{\top}\mathcal{D}^{2}_{f}(u,x)\xi+O(\theta^{3}).$ (3.1)
Here $D_{f}^{s}(u,x)$ is the matrix of the $s$th-order derivative of the function $f(u,v)$ with respect to $v$ at the point $x$ and for $x,y\in\mathbb{S}^{p-1}$, the vector $\xi$ and the angle $\theta\in(0,\pi)$ are defined by the tangent–normal decomposition
@@
Note that $\theta\in(0,\pi)$ is the angle between $x$ and $y$, and $\xi\in\mathbb{S}^{p-1}$ is a vector orthogonal to $x$, i.e., $\xi\in\Omega_{x}=\{\xi\in\mathbb{S}^{p-1}:\xi\perp x\}$. The intuition for (3.1) comes from the fact that $y-x\approx\theta\xi$ when $\theta$ approximates 0, which follows from a Taylor expansion of the function $\cos\theta$ and $\sin\theta$ at the point $\theta=0$. Moreover, the transformation of the area element of $\mathbb{S}^{p-1}$ corresponding to the tangent–normal decomposition (3.2) is given by
@@
see equation (2) in Marco Di Marzio and Taylor, (2014) or equation (1.5) in Hall et al., (1987). On the other hand, for a fixed $v\in\mathbb{S}^{p-1}$, we have the Taylor expansion for $f$
@@
where $\frac{\partial^{s}f(u,v)}{\partial u^{s}}$ denotes the partial derivative of order $s$ with respect to $u$.
Finally, we denote first and second derivative of the marginal density $f_{U}$ by $f_{U}^{\prime}$ and $f_{U}^{\prime\prime}$, respectively.
@@
$K(\cdot)$ is a symmetric density function supported on the interval (1,1)(-1,1), and $\int_{-1}^{1}K^{4}\left(u\right)\mathrm{d}u<\infty$; the kernel $L(\cdot)$ is the Langevin kernel defined in (2.10).
@@
The density $f$ of $Y$ is twice differentiable.
@@
    If the distribution of $Y$ is spherical symmetric, that is $\mathcal{M}^{2}=0$, we have
@@
    If the distribution of $Y$ is not spherical symmetric, that is $\mathcal{M}^{2}>0$, we have
@@
By Proposition 3.1, the statistic $\hat{\mathcal{M}}^{2}_{n}$ is in general not an unbiased estimator of $\mathcal{M}^{2}$. The bias of the kernel estimation with respect to the radius $\|Y\|$ is of order $h^{2}$ and it is of order $1/\kappa$ with respect to the direction $Y/\|Y\|$. Therefore, smaller values of $h$ and larger values of $\kappa$ are preferred to diminish the bias the estimator $\hat{\mathcal{M}}_{n}^{2}$. We now turn to the weak convergence of $\hat{\mathcal{M}}_{n}^{2}$. Throughout this paper the symbol $\stackrel{{\scriptstyle d}}{{\longrightarrow}}$ denotes weak convergence (convergence in distribution) of real valued random variables.
@@
If 3.1 and 3.2 are satisfied, $\kappa\to\infty$, $h\to 0$, $\kappa^{(p-1)/2}/(nh)\to 0$, and the distribution of $Y$ is spherical symmetric, that is $\mathcal{M}^{2}=0$, we have
@@
Note that the convergence rate in Theorem 3.1 is of order $\sqrt{n(n-1)h/{\kappa^{(p-1)/2}}}$, which is asymptotically greater than $\sqrt{n}$ as $\kappa^{(p-1)/2}/(nh)\to 0$, by assumption. On the other hand, weak convergence with a rate $\sqrt{n}$ is obtained in the case where the distribution of $Y$ is not spherical symmetric.
@@
Assume that 3.1 and 3.2 are satisfied and that $\kappa\to\infty$, $h\to 0$, $\kappa^{(p-1)/2}/(nh)\to 0$. If the distribution of $Y$ is not spherical symmetric, that is $\mathcal{M}^{2}>0$, we have
@@
Theorem 3.2 can be used to derive the weak convergence
$\sqrt{n}(\hat{\mathcal{M}}^{2}_{n}-{\mathcal{M}}^{2})\stackrel{{\scriptstyle d}}{{\longrightarrow}}{\cal N}(0,4\sigma^{2})$ (3.13)
in the case, where the distribution of $Y$ is not spherical symmetric, that is $\mathcal{M}^{2}>0$. However, some care is necessary when replacing $\mathbb{E}[\hat{\mathcal{M}}^{2}_{n}]$ by $\mathcal{M}^{2}$ as the error of this replacement has to be of order $o(1/\sqrt{n})$. By Proposition 3.1(ii) this yields to the condition
@@
This issue can be mitigated via a simple bias reduction using the statistic
@@
where $\hat{\mathcal{M}}^{2}_{n}(a\kappa)$ denotes the estimator $\hat{\mathcal{M}}^{2}_{n}$ using $a\kappa$ instead of $\kappa$ in the rapidly varying function $L(\cdot)$, see also Theorem 3 of Tsuruta, (2024). In addition, if we replace the kerne $K$ by its jackknife correction $\tilde{K}(x)=2\sqrt{2}K(\sqrt{2}x)-K(x)$, which satisfies $\phi_{2}(\tilde{K})=0$, we obtain from Proposition 3.1(ii)
@@
and (3.13) holds under the conditions $\sqrt{n}/\kappa^{3/2}+\sqrt{n}h^{3}\to 0$ and $\kappa^{(p-1)/2}/(nh)\to 0$, which can be satisfied, whenever $p\leq 6$. If $p>6$ corresponding results hold under additional smoothness assumptions using the higher order expansions
@@
where $u^{\otimes s}$ denotes its $s$-th Kroneckerian power of the vector $u$. For example, if $q=3$, it follows observing $\int_{\Omega_{x}}\xi^{\top}\xi^{\otimes 2}\omega_{p-2}(\xi)=0$, and $\int x^{3}K(x)dx=0$ ($K$ is symmetry by Assumption 3.1) that the rate in (3.16) can be improved to $O(\kappa^{-2}+h^{4})$.

### v2

@@ Abstract
Most of the work on checking spherical symmetry assumptions on the distribution of a $p$-dimensional random vector $Y$ focuses on testing for the null hypothesis of exact spherical symmetry. In this paper, we take a different point of view and consider the problem of testing if the deviation from spherical symmetry exceeds a given threshold. For this purpose, we propose a novel measure for the deviation from spherical symmetry, which is based on the minimum distance between the distribution of the vector $\big(|Y|,Y/|Y|)^{\top}$ and its best approximation by a distribution of a vector $\big(|Y_{s}|,Y_{s}/|Y_{s}|)^{\top}$ , where $Y_{s}$ follows a spherical distribution. We develop estimators for the minimum distance with corresponding statistical guarantees (provided by asymptotic theory) and use these results to develop statistical inference. Our approach assumes neither Gaussianity nor the existence of moments, and we demonstrate the applicability of our approach by means of a simulation study and a real data example.
@@ 1 Introduction
Spherical contoured models include multivariate Gaussian distribution, multivariate $t$- distribution and multivariate stable distribution find their applications in vector cardiology (mardia1999directional), portfolio theory (gupta2013elliptically), Mahalanobis distances (luis2005). There has been a wide interest in testing for the spherical symmetry, or a wider class of elliptical symmetry, see for example Smith01011977, ludwig1991, FANG199334, KOLTCHINSKII1998228, SCHOTT2002, HP2007, liangetal2008, PG2008, Henzeetal2014, holzman2020, babic2021, huang2023multivariate, banerjee2024consistent among others. A common feature of most of the cited references is that the authors use different approaches to construct tests for the null hypothesis of exact spherical symmetry. However, this assumption is rarely met because one does not believe that it is exactly (with mathematical equality) satisfied, but with the hope that the deviation from the null hypothesis is small such that (optimal) inference under the assumption of spherical symmetry is still reliable. Thus, strictly speaking, there are many situations where tests for exact sphericity are performed, although it is clear that this hypothesis is at most “approximately” satisfied.
@@
In this paper, we take a different point of view on the problem of validating the assumption of spherical symmetry on the distribution of a random vector $Y$. Instead of testing for exact spherical symmetry, we propose a measure of deviation from spherical symmetry. This measure is based on the minimum distance between the distribution of the vector $\big(\|Y\|,Y/\|Y\|)^{\top}$ and its best approximation by a distribution of a vector $\big(\|Y_{s}\|,Y_{s}/\|Y_{s}\|)^{\top}$ corresponding to a vector $Y_{s}$ with a spherical distribution, where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{p}$ (note that a random variable $Y_{s}$ has a spherical distribution if and only if $\|Y_{s}\|$ and $Y_{s}/\|Y_{s}\|$ are independent and $Y_{s}/\|Y_{s}\|$ is uniformly distributed on the $p$-dimensional sphere $\mathbb{S}_{p-1}$). In Section 2 we derive an explicit expression for this minimum distance, which is essentially the difference between the (squared) $L^{2}$-norm of the density of the vector $\big(\|Y_{s}\|,Y_{s}/\|Y_{s}\|)^{\top}$ and the (squared) $L^{2}$-norm of the density of the random variable $\|Y_{s}\|$ multiplied with the surface area of $\mathbb{S}^{p-1}$. Consequently, the estimation problem of the minimum distance boils down to estimating the squared $L^{2}$-norm of the density of a distribution on the sphere and a distribution on $\mathbb{R}^{+}$. While the problem of estimating the integrated squared density on $\mathbb{R}^{p}$ has been extensively studied from various perspectives including optimality and adaptivity (see hall1987estimation; bickel1988estimating; LaurantMassart2000; GineNickl2008, among many others), the corresponding problem on the sphere have not found much attention so far. We construct an estimator for the minimum distance from a sequence of independent identically distributed observations combining an estimator for the integrated squared density of the joint distribution of $Y/\|Y\|$ and $\|Y\|$ with an estimator for the integrated squared density of $\|Y\|$ on the non-negative line. We prove asymptotic normality of an appropriately centered and normalized version. Interestingly, from the theoretical perspective, there appear to be different scalings, depending on whether the distribution is spherically symmetric or not. In the latter case, the difference of the estimators converges at a rate $\sqrt{n}$ as in the case estimating the integral of the squared density (see, for example, bickel1988estimating). On the other hand, if the distribution of $Y$ is in fact spherically symmetric, the estimator converges at a rate $\sqrt{n(n-1)h/{\kappa^{(p-1)/2}}}$, where $h\to 0$ and $\kappa\to\infty$ denote smoothing parameters required for the estimation of the integrated squared densities of $\|Y\|$ and $Y/\|Y\|$, respectively.
@@
Based on these results, we develop several statistical applications. In Section 3.2 we derive an asymptotic confidence interval for the deviation of the distribution from sphericity. Moreover, we also construct tests for the hypothesis that the deviation from sphericity exceeds a given threshold. For example, if ${\cal M}^{2}$ denotes the measure of deviation we derive a consistent and asymptotic level $\alpha$-test for the hypotheses $H_{0}^{\rm eq}:{\cal M}^{2}>\Delta$ versus $H_{1}^{\rm eq}:{\cal M}^{2}\leq\Delta$, which allows to decide for approximate sphericity at a controlled type $I$ error. These procedures require the estimation of a complicated asymptotic variance, which we address by a Jackknife procedure. Moreover, as an alternative, we also develop in Section 4 pivotal inference for the measure ${\cal M}^{2}$, which does not require any estimation of asymptotic variances. Our approach is based on the self-normalization principle (see shao2015), which cannot be used directly in the present context. In particular, we have to derive the weak convergence of a process of non-parametric U-statistics with random elements on the sphere, and adapt it to obtain a pivotal statistic for the estimator of ${\cal M}^{2}$. Finally, we demonstrate the applicability of our approach through a small simulation study and a real data example.
@@ 2 A measure of deviation and its estimation
for some function $g$. Consequently, observing (2.1), it follows that $Y$ has a spherically symmetric distribution if and only if $U=\|Y\|$ and $V=Y/\|Y\|$ are independent and $V=Y/\|Y\|$ is uniformly distributed on the sphere with (constant) density
@@
Let $Y_{1},\ldots,Y_{n}$ denote independent identically distributed $p$-dimensional random variables and define $U_{i}=\|Y_{i}\|$; $V_{i}=Y_{i}/\|Y_{i}\|$ ($i=1,\ldots,n)$. By Proposition 2.1 the problem of estimating the minimum distance ${\cal M}^{2}$ boils down to the estimation of the integrated squared densities of the random variables $(U_{i},V_{i})$ and $U_{i}$. The latter can be estimated by standard methods (see, for example hall1987estimation; bickel1988estimating, among many others) and we use
@@
as estimate for the integrated squared density of $U$, where $h$ denotes a bandwidth, $K$ is a symmetric kernel function supported on (1,1)(-1,1) with integral $1$, for example the Epanechnikov kernel. The estimation of the first integral in (2.5) is more intricate as it corresponds to the (squared) $L^{2}$-norm of a density on $\mathbb{R}^{+}\times\mathbb{S}^{p-1}$. For this purpose, we use classical results from density estimation on the sphere, such as HallWatsonCabrera1987, and consider a rapidly varying function $L:\mathbb{R}^{p}\to\mathbb{R}$, which satisfies (i) $L$ is nonnegative and nondecreasing, (ii) for each $0<r<1$,
@@
$\hat{\cal M}_{n}^{(2)}:=\frac{1}{n(n-1)c_{1}(\kappa)h}\sum_{i\neq j}K\left(\frac{U_{i}-U_{j}}{h}\right)L(\kappa V_{i}^{\top}V_{j}),$ (2.8)
where $K(\cdot)$ and $h$ are the same kernel and bandwidth as used in $\hat{\cal M}_{n}^{(1)}$, $\kappa\to\infty$ is another bandwidth corresponding to the spherical part of the vector $(U,V)$, and the constant $c_{1}(\kappa)$ is given by
@@
for the kernel function $L(\cdot)$, where $\mathcal{I}_{\nu}$ is the modified Bessel function of the first kind (see also 9.6.18 on Page 376 of abramowitz1968handbook) of order $\nu$, i.e.,
@@ 3.1 Asymptotic theory
For the investigation of the bias of $\hat{\mathcal{M}}_{n}^{2}$, we recall a Taylor expansion for a twice continuously differentiable function $f:\mathbb{R}^{+}\times\mathbb{S}^{p-1}\to\mathbb{R}$ with respect to the argument $v$ on the sphere $\mathbb{S}^{p-1}$. According to equation (3) in DiMarzio03042014, we have
$f(u,y)=f(u,x)+\theta\xi^{\top}\mathcal{D}_{f}(u,x)+\frac{\theta^{2}}{2}\xi^{\top}\mathcal{D}^{2}_{f}(u,x)\xi+O(\theta^{3}),$ (3.1)
where $\mathcal{D}_{f}^{s}(u,x)$ is the matrix of the $s$th-order derivative of the function $f(u,v)$ with respect to $v$ at the point $x$ and for $x,y\in\mathbb{S}^{p-1}$, the vector $\xi$ and the angle $\theta\in(0,\pi)$ are defined by the tangent–normal decomposition
@@
Note that $\theta\in(0,\pi)$ is the angle between $x$ and $y$, and $\xi\in\mathbb{S}^{p-1}$ is a vector orthogonal to $x$, i.e., $\xi\in\Omega_{x}=\{z\in\mathbb{S}^{p-1}:z\perp x\}$. The intuition for (3.1) comes from the fact that $y-x\approx\theta\xi$ when $\theta$ approximates 0, which follows from a Taylor expansion of the function $\cos\theta$ and $\sin\theta$ at the point $\theta=0$. Moreover, the transformation of the area element of $\mathbb{S}^{p-1}$ corresponding to the tangent–normal decomposition (3.2) is given by
@@
see equation (2) in DiMarzio03042014 or equation (1.5) in HallWatsonCabrera1987. Meanwhile, for a fixed $v\in\mathbb{S}^{p-1}$, we have the Taylor expansion for $f$
@@
where $\partial^{s}f(u,v)/\partial u^{s}$ denotes the partial derivative of order $s$ with respect to $u$.
Finally, we denote first and second derivatives of the marginal density $f_{U}$ by $f_{U}^{\prime}$ and $f_{U}^{\prime\prime}$, respectively.
@@
$K(\cdot)$ is a symmetric function supported on the interval (1,1)(-1,1), $\int_{-1}^{1}K\left(u\right)=1$ and $\int_{-1}^{1}K^{4}\left(u\right)\mathrm{d}u<\infty$; the kernel $L(\cdot)$ is the Langevin kernel defined in (2.10).
@@
The density $f$ of $Y$ is twice continuously differentiable.
@@
    If the distribution of $Y$ is spherically symmetric, that is $\mathcal{M}^{2}=0$, we have
@@
    If the distribution of $Y$ is not spherically symmetric, that is $\mathcal{M}^{2}>0$, we have
@@
By Proposition 3.1, the statistic $\hat{\mathcal{M}}^{2}_{n}$ is an unbiased estimator of $\mathcal{M}^{2}$ only under spherical symmetry. When the exact spherical symmetry does not hold, the bias of the kernel estimation with respect to the radius $\|Y\|$ is of order $h^{2}$, and it is of order $1/\kappa$ with respect to the direction $Y/\|Y\|$. Therefore, smaller values of $h$ and larger values of $\kappa$ are preferred to diminish the bias of the estimator $\hat{\mathcal{M}}_{n}^{2}$. We now turn to the weak convergence of $\hat{\mathcal{M}}_{n}^{2}$. Throughout this paper, the symbol $\stackrel{{\scriptstyle d}}{{\longrightarrow}}$ denotes weak convergence (convergence in distribution) of real-valued random variables.
@@
If 3.1 and 3.2 are satisfied, $\kappa\to\infty$, $h\to 0$, $\kappa^{(p-1)/2}/(nh)\to 0$, and the distribution of $Y$ is spherically symmetric, that is $\mathcal{M}^{2}=0$, we have
@@
Note that the convergence rate in Theorem 3.1 is of order $\sqrt{n(n-1)h/{\kappa^{(p-1)/2}}}$, which is asymptotically greater than $\sqrt{n}$ as $\kappa^{(p-1)/2}/(nh)\to 0$, by assumption. On the other hand, weak convergence with a rate $\sqrt{n}$ is obtained in the case where the distribution of $Y$ is not spherically symmetric.
@@
Assume that 3.1 and 3.2 are satisfied and that $\kappa\to\infty$, $h\to 0$, $\kappa^{(p-1)/2}/(nh)\to 0$. If the distribution of $Y$ is not spherically symmetric, that is $\mathcal{M}^{2}>0$, we have
@@
Under exact sphericity (${\cal M}^{2}=0$), it follows from Theorem 3.1 that the variance of ${\cal M}^{2}_{n}$ is of order $\kappa^{(p-1)/2}/(n(n-1)h)$ and the estimator is consistent independently of the dimension.
In the case ${\cal M}^{2}>0$, the situation is more complicated. Theorem 3.2 can be used to derive the weak convergence
$\sqrt{n}(\hat{\mathcal{M}}^{2}_{n}-{\mathcal{M}}^{2})\stackrel{{\scriptstyle d}}{{\longrightarrow}}{\cal N}(0,4\sigma^{2}).$ (3.13)
However, some care is necessary when replacing $\mathbb{E}[\hat{\mathcal{M}}^{2}_{n}]$ by $\mathcal{M}^{2}$ as the error of this replacement has to be of order $o(1/\sqrt{n})$. By Proposition 3.1(ii) this yields to the condition
@@
This issue can be mitigated via a simple bias reduction
@@
where $\hat{\mathcal{M}}^{2}_{n}(a\kappa)$ denotes the estimator $\hat{\mathcal{M}}^{2}_{n}$ using $a\kappa$ instead of $\kappa$ in the rapidly varying function $L(\cdot)$ in (2.10), see also Theorem 3 of TSURUTA2024105338. In addition, if we replace the kernel $K(\cdot)$ by its Jackknife correction $\tilde{K}(x)=2\sqrt{2}K(\sqrt{2}x)-K(x)$, which satisfies $\phi_{2}(\tilde{K})=0$, we obtain from Proposition 3.1(ii)
@@
and (3.13) holds under the conditions $\sqrt{n}/\kappa^{3/2}+\sqrt{n}h^{3}\to 0$ and $\kappa^{(p-1)/2}/(nh)\to 0$, which can be satisfied, whenever $p\leq 6$, for example $\kappa=n^{\frac{3}{p+3}},h=n^{-\frac{3}{2(p+3)}}$. If $p>6$, corresponding results can be derived under additional smoothness assumptions using the higher order expansions
@@
where $u^{\otimes s}$ denotes the $s$-th Kroneckerian power of the vector $u$. For example, if $q=3$, it follows, observing $\int_{\Omega_{x}}\xi^{\top}\xi^{\otimes 2}\omega_{p-2}(\xi)=0$ and $\int x^{3}K(x)dx=0$ ($K$ is symmetric by Assumption 3.1), that the rate in (3.16) can be improved to $O(\kappa^{-2}+h^{4})$, and (3.13) holds under the conditions
$\sqrt{n}/\kappa^{2}+\sqrt{n}h^{4}\to 0,$ (3.19)
which can be satisfied, whenever $p\leq 8$, for example $\kappa=n^{\frac{3}{p+4}},h=n^{-\frac{3}{2(p+4)}}$.